The left ventricle of the heart accelerates blood from rest to a velocity of +24 cm/s.

(a) If the displacement of the blood during the acceleration is +2.3 cm, determine its acceleration (in cm/s2).


(b) How much time does blood take to reach its final velocity?
s

2. A car is traveling at 11.0 m/s, and the driver sees a traffic light turn red. After 0.570 s (the reaction time), the driver applies the brakes, and the car decelerates at 6.00 m/s2. What is the stopping distance of the car, as measured from the point where the driver first sees the red light?
Answer #1

a) Using equation of the motion V2 =U2 +2aS ....1

here U=0 (rest ) , a=? , S=2.3 cm , V=24 cm/s

put these value in equation 1

(24)2 =0+2a*2.3

a=242/2*2.3

a=125.22 cm/s2 Answer

b) Again Using equation of the motion v=u+at

t=v-u/a

t=(24-0)/125.22

t=0.192 s

2) Distance travel in the reaction time t=0.570 s d1=vt =11*0.57

d1 =6.27 m

distance travel during deceleration (a= -6 m/s2)

Using equation of the motion V2 =u2 +2ad2

d2 =V2 -u2 /2a

here V= 0 (stop) , u=11 m/s . a=-6 m/s2

d2 =0-112/(-2*6)

d2 =10.083 m

total distance =d1+d2 =6.27+10.083

D=16.353 m Answer